Solve the differential equation
dy/dx= x+2y-3/2x+y-3
1
Expert's answer
2020-12-08T07:46:26-0500

"\displaystyle

\frac{\mathrm{d}y}{\mathrm{d}x} = -\frac{x + 2y - 3}{2x + y - 3}\\


(2x + y - 3)\mathrm{d}y + (x + 2y - 3)\mathrm{d}x = 0\\


M\mathrm{d}y + N\mathrm{d}x = 0\\


M_x = \frac{\partial M}{\partial x} = 2\\


N_x = \frac{\partial N}{\partial y} = 2\\

\textsf{Since}\,\, M_y = N_x\,\,\textsf{the differential equation is exact}\\


\frac{\partial z}{\partial y} = 2x + y - 3\\

z = \int(2x + y - 3)\, \mathrm{d}y\\

z = 2xy + \frac{y^2}{2} - 3y + f(x)\\


\frac{\partial z}{\partial x} = x + 2y - 3\\

z = \int (x + 2y - 3)\,\mathrm{d}x\\

z = \frac{x^2}{2} + 2xy - 3x + f(y)\\


z = 2xy + \frac{x^2 + y^2}{2} - 3(x + y)"